THE SUM OF BOTH DIGITS, OF EITHER OF TWO TWO-DIGET NUMBERS, IN WHATEVER ORDER THE DIGETS ARE WRITTEN, IS 9. THE SQUARE OF EITHER OF THE DIGETS OF EITHER NUMBER, MINUS THE PRODUCT OF BOTH DIGITS, PLUS THE SQUARE OF THE OTHER DIGIT IS THE NUMBER 21. THE NUMBERS ARE:
A)36,63
B)81,18
C)27,72
D)45,54

Answer by edjones(8007)

4^2+5^2-20=16+25-20=21